OpenStudy (anonymous):

if f(x)= tan5x- sin2x, find f(pi/6)

OpenStudy (mertsj):

\[\tan \frac{5\pi}{6}-\sin \frac{\pi}{3}=\frac{-\sqrt{3}}{3}-\frac{\sqrt{3}}{2}=\frac{-2\sqrt{3}-3\sqrt{3}}{6}=\frac{-5\sqrt{3}}{6}\]

OpenStudy (anonymous):

thank you. :)

OpenStudy (mertsj):

yw